OpenStudy (ttp):

What is the value of the expression b times the quantity begin square root a squared minus nine end square root end quantity minus b squared when a equals five and b equals one? 0 3 4 twelve

OpenStudy (johnweldon1993):

\[\large b \sqrt{a^2 - 9} - b^2\] When a = 5 and b = 1 we have \[\large \sqrt{(5)^2 - 9} - 1\] \[\large \sqrt{25 - 9} - 1\] \[\large \sqrt{16} - 1\]etc.. until you get...?

OpenStudy (ttp):

I might neede help

OpenStudy (johnweldon1993):

Okay :) So we left off at \[\large \sqrt{16} - 1\] We know that the square root of 16 is 4...(since 4 x 4= 16) So we have \[\large 4 - 1 = ?\]

OpenStudy (ttp):

3

OpenStudy (johnweldon1993):

And there's you answer :)